OpenStudy (anonymous):

Find the values of 'x' satisfying this solution. [x] + 1/[2x] ={x} +1/3 where [ ] represents greatest integer function and { } represents fractional part. I want to know the procedure.

OpenStudy (anonymous):

somebody help please.........

OpenStudy (anonymous):

you can write \(x=n+\left\{ x \right\}\) where \(n=[x]\) and \(0\le\left\{ x \right\}<1\), put it in the equation:\[n+\frac{1}{[2(n+\left\{ x \right\})]}=\left\{ x \right\}+\frac{1}{3}\]you will get\[n+\frac{1}{2n+[2\left\{ x \right\}]}=\left\{ x \right\}+\frac{1}{3}\](why?)now note that \([2\left\{ x \right\}]\) can either be \(0\) or \(1\) (why?) case 1: \([2\left\{ x \right\}]=0\) \[n+\frac{1}{2n}=\left\{ x \right\}+\frac{1}{3}\]\(n\) is a positive integer, hence no answer from here (why?) case 2: \([2\left\{ x \right\}]=1\) \[n+\frac{1}{2n+1}=\left\{ x \right\}+\frac{1}{3}\]\(n\) is a positive integer, hence no answer from here (why?).